# and quantum information science

### v2

@@ 5.6.1 Free quantum field theories
The study of complexity in scalar bosonic QFT began with [99, 100]. In these papers, the authors considered the following Hamiltonian for a scalar QFT in $d$-dimensional spacetime:
@@
where $\hat{\rho}$ is the Gaussian density matrix, and $\Omega^{[ab]}$ encodes the anti-symmetric commutation relations of the phase space variables. The covariance matrices were not the original formulation used to study complexity in bosonic QFT. Originally, the studies only focused on squeezing operations, and therefore expressing everything in terms of the wavefunction was sufficient. Nevertheless the use of covariance matrices permits for more general transformations between Gaussian states which was employed in later studies, see e.g., [104, 105]. We will therefore stick with this approach in our presentation.
@@
When evaluating complexity in QFTs, there are several choices to be made. Initial studies [99, 100] selected cost functions according to equation (25), with $q_{I}=1$ and $p=1,2$, and only squeezing operations were considered. The target state was taken to be the regularized ground state of the bosonic field theory (40). Regularization was implemented either via a lattice discretization or a sharp momentum cutoff. The reference state was chosen as the ground state of the Hamiltonian
@@
The original papers [99, 100] focused primarily on the structure of divergences and found some degree of agreement with holographic results. In particular, they calculated the complexity of the vacuum state of a free bosonic theory, which reads:
$\begin{split}\mathcal{C}_{p=1}^{\text{upper bound}}&\ =\frac{1}{2}\int d^{d}k\left|\log(\omega_{k}/\mu)\right|\simeq\frac{\mathcal{V}}{2\delta^{d-1}}|\log(\tilde{\mu})|+\ldots,\\ \mathcal{C}_{p=2}&\ =\frac{1}{2}\sqrt{\int_{k}\left[\log(\omega_{k}/\mu)\right]^{2}}\simeq\frac{1}{2}\left(\frac{\mathcal{V}}{\delta^{d-1}}\right)^{\frac{1}{2}}|\log(\tilde{\mu})|+\ldots,\end{split}$ (46)
where for $p=1$ only an upper bound was found. In these equations $\mathcal{V}$ stands for the spatial volume of the system, $\delta$ is an ultraviolet cutoff (for example a lattice spacing) and the dots stand for less divergent terms (see the review [2]). These results should be contrasted with those obtained from holography (125). We will later see that the $\mathcal{C}_{p=1}$ expression matches well with the volume-law divergence found in holographic calculations, see Eq. (125). Additionally, the choice of reference scale in the field theory parallels the choice of the counterterm scale $L_{ct}$ in holography, as we will see in, e.g., Equation (128a).
Subsequent studies explored a variety of alternative cost functions, free theories with other fields such as charged bosons, fermions, and gauge fields, weakly interacting theories and other target states, including the thermofield-double state, and mixed states [111, 104, 105, 103, 106, 107, 125, 109, 110, 126, 127, 128, 129, 130]. These studies concluded that, in general, the structure of divergences observed in QFT complexity calculations qualitatively matches the results from holography. This is not particularly surprising, since at very short distances (or equivalently, near the cutoff scale), the precise details of the theory tend to become irrelevant. On the other hand, the time dynamics of the theory shows significant deviations from holographic results, which is to be expected, as free and chaotic theories typically exhibit very different dynamical behavior. In particular, a natural expectation is that in free systems of finite volume, after including penalties on Gaussian gates coupling distant parts of the system, complexity growth associated with Hamiltonian time evolution will terminate at times of the order of the volume (and even sooner without the penalties [105]). In chaotic systems, we expect complexity growth to occur over significantly longer timescales, exponential in the volume.
@@ 5.6.2 Two dimensions - complexity of Virasoro circuits
Such challenges forced the community to approach the problem of Nielsen’s complexity in QFT in a creative way. Progress can be traced back to [40], whose authors proposed to study Nielsen’s complexity for conformal transformations of 2-dimensional CFTs, i.e., coordinate transformations that preserve the form of the metric up to an overall (coordinate-dependent) scalar factor.
@@
In more detail, [40] and subsequent studies focused on two-dimensional CFTs living on the Lorentzian cylinder. As illustrated in Fig. 6, we parametrize the spatial circle of unit radius by $\sigma$ and the physical time by $t$. In two dimensions, the energy-momentum tensor of a CFT has two independent components that commute with each other. In the following, we will predominantly focus on one of them only, $T(\sigma)\equiv\frac{1}{2}T_{tt}+\frac{1}{2}T_{t\sigma}$. The algebraic structure underlying this subsection’s construction is the Virasoro algebra
$[L_{n},L_{m}]=(n-m)L_{n+m}+\frac{c}{12}n(n^{2}-1)\delta_{n+m,0}\,,$ (47)
@@
Figure 6: Two kinds of circuits that can be considered in the context of conformal transformations, or, more generally, QFT (in this latter scenario, one should think of $\sigma$ as a set of coordinates on spatial slices). Case (a): The circuit parameter $\tau$ is an auxiliary parameter whose increase takes one CFT state to another CFT state. The transformation does not lead to a shift in the physical time $t$. This is the setting addressed in [40]. Case (b): The circuit parameter $\tau$ is identified with the physical time $t$, as considered for the first time in [131]. Figure adapted from [131].
The setup of interest for [40], as well as subsequent works including [132, 133, 133, 134], is given by continuous unitary circuits
[…]
@@
In considering the Nielsen’s complexity, there is an important choice to be made regarding the interpretation of the $\tau$ parameter [131]. The first choice is to regard it as an auxiliary parameter in relation to the physical time $t$, see case (a) in Fig. 6. For this choice, relevant for [40], the function $\epsilon(\tau)$ associated with the diffeomorphism $\sigma\rightarrow f(\tau,\sigma)$ is given by
@@
The papers [40, 132, 133, 135] uncovered and developed a very beautiful relation between optimizing cost functions of the form
[…]
@@
with $a$ and $b$ constant and $\delta(\sigma-\kappa)$ the Dirac delta function, the solutions of the optimization problem are the solutions of the following paradigmatic PDEs [133]
@@
Furthermore, Refs. [133, 135] considered the Fubini-Study cost function, which corresponds to Eq. (53) with
@@
and solved the associated optimization perturbatively for sample conformal transformations close to identity. This result was subsequently used for comparisons with holography, building on [136], see Sec. 7.
There are further generalizations of the approach described above. Reference [132] connected with the rich literature of the Kac-Moody groups by incorporating an additional global symmetry. Furthermore, in Ref. [137] a scalar primary operator was added as an additional generator on top of the energy-momentum tensor operator. The main limitation of the works prior to [137] is that the previously described unitary circuits remain bounded inside a conformal family (the so-called Verma module) of the Virasoro algebra. Results in [137] concern trajectories that move between different Verma modules in a CFT, revealing the dependence of the Fubini-Study cost of such circuits on the source function of the primary generator. This investigation constitutes an important step towards arriving at a general picture of Nielsen’s complexity with local operators viewed as generators, as advocated above.
Finally, another line of works devoted to the study of complexity in CFT by examining trajectories in which the physical time $t$ is the circuit parameter $\tau$ [131, 44, 134], see Fig. 6 case (b). In the context of the examples considered, this approach requires slightly different choices of the circuit velocity $\epsilon(t,\sigma)$. The main advantage of this approach is that it makes direct contact with the spacetime (path integral) formulation of QFTs and with holography [44].
@@ 5.6.3 Circuits in the conformal group in $d\geq 2$
This framework is similar to the qubit setup in Sec. 5.1, with the exception that we now consider a unitary representation of a (possibly non-compact) Lie group $G$ [^18] One can always build a unitary and finite-dimensional representation of a Lie group by using Euclidean generators. For concreteness, a generic trajectory on the conformal group can be constructed out of the generators $P_{\mu}$, $D$, $L_{\mu\nu}$ and $K_{\mu}$ (which generate translations, dilatations, rotations and special conformal transformations, respectively) as follows,
@@
Let us discuss Nielsen’s state complexity associated with a target state $|\psi_{\rm R}\rangle$, reached by acting on a primary state $|\Delta\rangle$ with a unitary operator of the form (56). The projection of a Riemannian right-invariant cost function from a Lie group $G$ to the coset space $G/H$ (where $H$ is the maximal proper subgroup of a state) is unique, and it is defined by a pseudo-Riemannian submersion [138, 139]. This provides a systematic procedure for inducing a metric from the conformal group $\mathrm{SO}(d,2)$ to the coset space $\mathcal{B}=\frac{\mathrm{SO}(d,2)}{\mathrm{SO}(2)\times\mathrm{SO}(d)}$. Moreover, the projection procedure is equivalent to two other ways of generating a metric over the quotient: the minimization in Eq. (23), and the use of a geometric action associated with the geometry of coadjoint orbits [140]. A coadjoint orbit is the space of all equivalent configurations that can be reached by acting with the symmetry group on a point in the dual space of the corresponding Lie algebra. Physically, it represents a phase space corresponding to a fixed value of a conserved charge, where the symmetry group of the system allows movement within this space. Two cost function candidates were considered in the literature in this context [42]
@@
Both cost functions have some limitations because of the absence of penalty factors associated with different gates. Nevertheless, they provide a good starting point. In particular, the Fubini-Study (FS) line element $ds_{\rm FS}$ given by Eq. (57b) is the natural metric on the coset space $\mathcal{B}$ of the conformal group, whose geodesics are known. The cost function in Eq. (57b) can be related to a Kähler metric associated with coadjoint orbits in the dual space [42]. It is still an open problem to generalize the results of [40, 42] to account for non-trivial penalty factors; this is the subject of ongoing research [140]. Finally, Ref. [42] found an explicit form of the complexity metric for higher-dimensional CFTs. In terms of the control functions in Eq. (56), their result has a precise analogue in terms of distances between geodesics in AdS, as we will discuss in Sec. 7.6.
@@ 6 Paradigms for complexity II: Krylov and spread complexities
One approach in quantum mechanics is to quantify the growth of different operators over time via the notion of Krylov complexity developed in [21]. Alternatively, we could ask how different initial states dynamically explore the Hilbert space during unitary evolution via the notion of spread complexity developed in [22]. These approaches are related, but each of them is convenient for answering different questions. In both approaches, the key insight is to recognize that, while physicists often analyze quantum systems by diagonalizing the Hamiltonian to extract the density of states, it is much more convenient to tri-diagonalize the Hamiltonian if we want to study dynamics. The diagonal and immediately off-diagonal components of the triadiagonalized Hamiltonian are called the Lanczos coefficients, and any quantum dynamics can be written as a one-dimensional hopping chain in terms of the Lanczos coefficients and the associated Krylov basis for the Hilbert space. Among other things, this representation of a quantum system makes it possible to explicitly quantify and compute how a quantum wavefunction spreads under Hamiltonian dynamics across the Hilbert space, or how an operator spreads in operator space.
Below, we review the background for Krylov and spread complexities, and explain how to compute these quantities. We present numerical methods, analytical results in systems with a lot of symmetry, and a general analytical formula relating the density of states of a system to its Lanczos coefficients. We also present a formula for the correlations in the Lanczos coefficients of Random Matrix Theories. Along the way, we relate these ideas to Nielsen complexity, and describe applications to the study of chaos and integrability. In the next Section, devoted to holography, we will describe a precise relationship between the length of wormholes in the two-dimensional JT gravity, and the spread complexity of the Thermofield Double state of the dual SYK model, see Sec. 7.5.1. Readers may also wish to consult the excellent recent review [3].
@@ 6.1 Spread complexity of evolving states
where $H$ is a time-independent Hamiltonian. Since the evolution is unitary, the state at any given time is just a fixed vector in the Hilbert space. Standard methods, including the Out-Of-Time-Order-Correlation (OTOC, see definitions in Sec. 2), quantify how time evolution spreads a collection of nearby states through the Hilbert space. At least at early times, this process can be characterized by computing quantities akin to the Lyapunov exponent. Instead, we will quantify how widely the trajectory of a single state explores the Hilbert space over time. One way of doing this is to measure the spread of the support of a time-evolving state in some fixed basis. Which basis should we pick? The early work of Kolmogorov, defining the complexity of sequences as the size of the smallest Turing machine that produces them, suggested a paradigm: the complexity of a system should be measured in terms of the size of its minimal representation. Taking this perspective, we could define the complexity of an evolving state by minimizing the spread of the wavefunction over all possible bases. The authors of [22] proved that there is a unique orthonormal minimizing basis throughout a finite time interval for continuous evolution, and at all times in the case of discrete evolution. This is the Krylov basis which can be constructed by a variety of methods described below.
For discrete time evolution, a simple argument explains how the Krylov basis minimizes the spread of the wavefunction. Consider an initial state $\lvert\psi_{0}\rangle$, and a basis supporting the state on a single basis element $\lvert K_{0}\rangle=\lvert\psi_{0}\rangle$. Time evolution produces the state at time 1 as $\lvert\psi_{1}\rangle=U_{1}\lvert\psi_{0}\rangle$, where $U_{1}$ is a unitary operator. By picking the second basis element $\lvert K_{1}\rangle$ to be proportional to the part of $\lvert\psi_{1}\rangle$ orthogonal to $\lvert\psi_{0}\rangle$, we can support $\lvert\psi_{1}\rangle$ on two basis elements, $\{\lvert K_{0}\rangle,\lvert K_{1}\rangle\}$. Continuing in this way by recursively orthogonalizing the time-evolving state, we produce a basis that obviously minimizes the support of the state at any time. The analogous proof for continuous time evolution is more subtle and is described in [22].

### v3

@@ 5.6.1 Free quantum field theories
The study of complexity in scalar bosonic QFT began with Jefferson:2017sdb, Chapman:2017rqy. In these papers, the authors considered the following Hamiltonian for a scalar QFT in $d$-dimensional spacetime:
@@
where $\hat{\rho}$ is the Gaussian density matrix, and $\Omega^{[ab]}$ encodes the anti-symmetric commutation relations of the phase space variables. The covariance matrices were not the original formulation used to study complexity in bosonic QFT. Originally, the studies only focused on squeezing operations, and therefore expressing everything in terms of the wavefunction was sufficient. Nevertheless the use of covariance matrices permits for more general transformations between Gaussian states which was employed in later studies, see e.g., Hackl:2018ptj, Chapman:2018hou. We will therefore stick with this approach in our presentation.
@@
When evaluating complexity in QFTs, there are several choices to be made. Initial studies Jefferson:2017sdb, Chapman:2017rqy selected cost functions according to equation (25), with $q_{I}=1$ and $p=1,2$, and only squeezing operations were considered. The target state was taken to be the regularized ground state of the bosonic field theory (40). Regularization was implemented either via a lattice discretization or a sharp momentum cutoff. The reference state was chosen as the ground state of the Hamiltonian
@@
The original papers Jefferson:2017sdb, Chapman:2017rqy focused primarily on the structure of divergences and found some degree of agreement with holographic results. In particular, they calculated the complexity of the vacuum state of a free bosonic theory, which reads:
[…]
where for $p=1$ only an upper bound was found. In these equations $\mathcal{V}$ stands for the spatial volume of the system, $\delta$ is an ultraviolet cutoff (for example a lattice spacing) and the dots stand for less divergent terms (see the review Chapman:2021jbh). These results should be contrasted with those obtained from holography (125). We will later see that the $\mathcal{C}_{p=1}$ expression matches well with the volume-law divergence found in holographic calculations, see Eq. (125). Additionally, the choice of reference scale in the field theory parallels the choice of the counterterm scale $L_{ct}$ in holography, as we will see in, e.g., Equation (128a).
Subsequent studies explored a variety of alternative cost functions, free theories with other fields such as charged bosons, fermions, and gauge fields, weakly interacting theories and other target states, including the thermofield-double state, and mixed states Balasubramanian:2018hsu, Hackl:2018ptj, Chapman:2018hou, Khan:2018rzm, Bhattacharyya:2018bbv, Guo:2018kzl, Cotler:2018ufx, Bernamonti:2019zyy, Caceres:2019pgf, Camargo:2020yfv, Hashimoto:2017fga, Chapman:2019clq, Ge:2019mjt, Camargo:2018eof. These studies concluded that, in general, the structure of divergences observed in QFT complexity calculations qualitatively matches the results from holography. This is not particularly surprising, since at very short distances (or equivalently, near the cutoff scale), the precise details of the theory tend to become irrelevant. On the other hand, the time dynamics of the theory shows significant deviations from holographic results, which is to be expected, as free and chaotic theories typically exhibit very different dynamical behavior. In particular, a natural expectation is that in free systems of finite volume, after including penalties on Gaussian gates coupling distant parts of the system, complexity growth associated with Hamiltonian time evolution will terminate at times of the order of the volume (and even sooner without the penalties Chapman:2018hou). In chaotic systems, we expect complexity growth to occur over significantly longer timescales, exponential in the volume.
@@ 5.6.2 Two dimensions - complexity of Virasoro circuits
Such challenges forced the community to approach the problem of Nielsen’s complexity in QFT in a creative way. Progress can be traced back to Caputa:2018kdj, whose authors proposed to study Nielsen’s complexity for conformal transformations of 2-dimensional CFTs, i.e., coordinate transformations that preserve the form of the metric up to an overall (coordinate-dependent) scalar factor.
@@
In more detail, Caputa:2018kdj and subsequent studies focused on two-dimensional CFTs living on the Lorentzian cylinder. As illustrated in Fig. 6, we parametrize the spatial circle of unit radius by $\sigma$ and the physical time by $t$. In two dimensions, the energy-momentum tensor of a CFT has two independent components that commute with each other. In the following, we will predominantly focus on one of them only, $T(\sigma)\equiv\frac{1}{2}T_{tt}+\frac{1}{2}T_{t\sigma}$. The algebraic structure underlying this subsection’s construction is the Virasoro algebra
$L_{n},L_{m}]=(n-m)L_{n+m}+\frac{c}{12}n(n^{2}-1)\delta_{n+m,0}\,$, (47)
@@
Figure 6: Two kinds of circuits that can be considered in the context of conformal transformations, or, more generally, QFT (in this latter scenario, one should think of $\sigma$ as a set of coordinates on spatial slices). Case (a): The circuit parameter $\tau$ is an auxiliary parameter whose increase takes one CFT state to another CFT state. The transformation does not lead to a shift in the physical time $t$. This is the setting addressed in Caputa:2018kdj. Case (b): The circuit parameter $\tau$ is identified with the physical time $t$, as considered for the first time in Erdmenger:2021wzc. Figure adapted from Erdmenger:2021wzc.
The setup of interest for Caputa:2018kdj, as well as subsequent works including Erdmenger:2020sup, Flory:2020eot, Flory:2020eot, deBoer:2023lrd, is given by continuous unitary circuits
$U(\tau)=\reflectbox{$\vec{\reflectbox{$\mathcal{P}$}}$}\exp\left[\int^{\tau}_{0}Q(\tau^{\prime})d\tau^{\prime}\right]\,$, (49)
@@
In considering the Nielsen’s complexity, there is an important choice to be made regarding the interpretation of the $\tau$ parameter Erdmenger:2021wzc. The first choice is to regard it as an auxiliary parameter in relation to the physical time $t$, see case (a) in Fig. 6. For this choice, relevant for Caputa:2018kdj, the function $\epsilon(\tau)$ associated with the diffeomorphism $\sigma\rightarrow f(\tau,\sigma)$ is given by
@@
The papers Caputa:2018kdj, Erdmenger:2020sup, Flory:2020eot, Flory:2020dja uncovered and developed a very beautiful relation between optimizing cost functions of the form
$F[\epsilon]=\sqrt{\int_{0}^{2\pi}d\sigma\int_{0}^{2\pi}d\kappa\,\Pi(\sigma-\kappa)\,\epsilon(\tau,\sigma)\,\epsilon(\tau,\kappa)}\,$, (53)
@@
with $a$ and $b$ constant and $\delta(\sigma-\kappa)$ the Dirac delta function, the solutions of the optimization problem are the solutions of the following paradigmatic PDEs Flory:2020eot
@@
Furthermore, Refs. Flory:2020eot, Flory:2020dja considered the Fubini-Study cost function, which corresponds to Eq. (53) with
@@
and solved the associated optimization perturbatively for sample conformal transformations close to identity. This result was subsequently used for comparisons with holography, building on Flory:2018akz, see Sec. 7.
There are further generalizations of the approach described above. Reference Erdmenger:2020sup connected with the rich literature of the Kac-Moody groups by incorporating an additional global symmetry. Furthermore, in Ref. Erdmenger:2024xmj a scalar primary operator was added as an additional generator on top of the energy-momentum tensor operator. The main limitation of the works prior to Erdmenger:2024xmj is that the previously described unitary circuits remain bounded inside a conformal family (the so-called Verma module) of the Virasoro algebra. Results in Erdmenger:2024xmj concern trajectories that move between different Verma modules in a CFT, revealing the dependence of the Fubini-Study cost of such circuits on the source function of the primary generator. This investigation constitutes an important step towards arriving at a general picture of Nielsen’s complexity with local operators viewed as generators, as advocated above.
Finally, another line of works devoted to the study of complexity in CFT by examining trajectories in which the physical time $t$ is the circuit parameter $\tau$ Erdmenger:2021wzc, Erdmenger:2022lov, deBoer:2023lrd, see Fig. 6 case (b). In the context of the examples considered, this approach requires slightly different choices of the circuit velocity $\epsilon(t,\sigma)$. The main advantage of this approach is that it makes direct contact with the spacetime (path integral) formulation of QFTs and with holography Erdmenger:2022lov.
@@ 5.6.3 Circuits in the conformal group in $d\geq 2$
This framework is similar to the qubit setup in Sec. 5.1, with the exception that we now consider a unitary representation of a (possibly non-compact) Lie group GG. [^21] One can always build a unitary and finite-dimensional representation of a Lie group by using Euclidean generators. For concreteness, a generic trajectory on the conformal group can be constructed out of the generators $P_{\mu}$, $D$, $L_{\mu\nu}$ and $K_{\mu}$ (which generate translations, dilatations, rotations and special conformal transformations, respectively) as follows,
@@
Let us discuss Nielsen’s state complexity associated with a target state $|\psi_{\rm R}\rangle$, reached by acting on a primary state $|\Delta\rangle$ with a unitary operator of the form (56). The projection of a Riemannian right-invariant cost function from a Lie group $G$ to the coset space $G/H$ (where $H$ is the maximal proper subgroup of a state) is unique, and it is defined by a pseudo-Riemannian submersion oneill1983semiriemannian, Besse:1987pua. This provides a systematic procedure for inducing a metric from the conformal group $\mathrm{SO}(d,2)$ to the coset space $\mathcal{B}=\frac{\mathrm{SO}(d,2)}{\mathrm{SO}(2)\times\mathrm{SO}(d)}$. Moreover, the projection procedure is equivalent to two other ways of generating a metric over the quotient: the minimization in Eq. (23), and the use of a geometric action associated with the geometry of coadjoint orbits CFTpen. A coadjoint orbit is the space of all equivalent configurations that can be reached by acting with the symmetry group on a point in the dual space of the corresponding Lie algebra. Physically, it represents a phase space corresponding to a fixed value of a conserved charge, where the symmetry group of the system allows movement within this space. Two cost function candidates were considered in the literature in this context Chagnet:2021uvi
@@
Both cost functions have some limitations because of the absence of penalty factors associated with different gates. Nevertheless, they provide a good starting point. In particular, the Fubini-Study (FS) line element $ds_{\rm FS}$ given by Eq. (57b) is the natural metric on the coset space $\mathcal{B}$ of the conformal group, whose geodesics are known. The cost function in Eq. (57b) can be related to a Kähler metric associated with coadjoint orbits in the dual space Chagnet:2021uvi. It is still an open problem to generalize the results of Caputa:2018kdj, Chagnet:2021uvi to account for non-trivial penalty factors; this is the subject of ongoing research CFTpen. Finally, Ref. Chagnet:2021uvi found an explicit form of the complexity metric for higher-dimensional CFTs. In terms of the control functions in Eq. (56), their result has a precise analogue in terms of distances between geodesics in AdS, as we will discuss in Sec. 7.6.
@@ 6 Paradigms for complexity II: Krylov and spread complexities
One approach in quantum mechanics is to quantify the growth of different operators over time via the notion of Krylov complexity developed in Parker:2018yvk. Alternatively, we could ask how different initial states dynamically explore the Hilbert space during unitary evolution via the notion of spread complexity developed in Balasubramanian:2022tpr. These approaches are related, but each of them is convenient for answering different questions. In both approaches, the key insight is to recognize that, while physicists often analyze quantum systems by diagonalizing the Hamiltonian to extract the density of states, it is much more convenient to tri-diagonalize the Hamiltonian if we want to study dynamics. The diagonal and immediately off-diagonal components of the triadiagonalized Hamiltonian are called the Lanczos coefficients, and any quantum dynamics can be written as a one-dimensional hopping chain in terms of the Lanczos coefficients and the associated Krylov basis for the Hilbert space. Among other things, this representation of a quantum system makes it possible to explicitly quantify and compute how a quantum wavefunction spreads under Hamiltonian dynamics across the Hilbert space, or how an operator spreads in operator space.
Below, we review the background for Krylov and spread complexities, and explain how to compute these quantities. We present numerical methods, analytical results in systems with a lot of symmetry, and a general analytical formula relating the density of states of a system to its Lanczos coefficients. We also present a formula for the correlations in the Lanczos coefficients of Random Matrix Theories. Along the way, we relate these ideas to Nielsen complexity, and describe applications to the study of chaos and integrability. In the next Section, devoted to holography, we will describe a precise relationship between the length of wormholes in the two-dimensional JT gravity, and the spread complexity of the Thermofield Double state of the dual SYK model, see Sec. 7.5.1. Readers may also wish to consult the excellent recent review Nandy:2024htc.
@@ 6.1 Spread complexity of evolving states
where $H$ is a time-independent Hamiltonian. Since the evolution is unitary, the state at any given time is just a fixed vector in the Hilbert space. Standard methods, including the Out-Of-Time-Order-Correlation (OTOC, see definitions in Sec. 2), quantify how time evolution spreads a collection of nearby states through the Hilbert space. At least at early times, this process can be characterized by computing quantities akin to the Lyapunov exponent. Instead, we will quantify how widely the trajectory of a single state explores the Hilbert space over time. One way of doing this is to measure the spread of the support of a time-evolving state in some fixed basis. Which basis should we pick? The early work of Kolmogorov, defining the complexity of sequences as the size of the smallest Turing machine that produces them, suggested a paradigm: the complexity of a system should be measured in terms of the size of its minimal representation. Taking this perspective, we could define the complexity of an evolving state by minimizing the spread of the wavefunction over all possible bases. The authors of Balasubramanian:2022tpr proved that there is a unique orthonormal minimizing basis throughout a finite time interval for continuous evolution, and at all times in the case of discrete evolution. This is the Krylov basis which can be constructed by a variety of methods described below.
For discrete time evolution, a simple argument explains how the Krylov basis minimizes the spread of the wavefunction. Consider an initial state $\lvert\psi_{0}\rangle$, and a basis supporting the state on a single basis element $\lvert K_{0}\rangle=\lvert\psi_{0}\rangle$. Time evolution produces the state at time 1 as $\lvert\psi_{1}\rangle=U_{1}\lvert\psi_{0}\rangle$, where $U_{1}$ is a unitary operator. By picking the second basis element $\lvert K_{1}\rangle$ to be proportional to the part of $\lvert\psi_{1}\rangle$ orthogonal to $\lvert\psi_{0}\rangle$, we can support $\lvert\psi_{1}\rangle$ on two basis elements, $\{\lvert K_{0}\rangle,\lvert K_{1}\rangle\}$. Continuing in this way by recursively orthogonalizing the time-evolving state, we produce a basis that obviously minimizes the support of the state at any time. The analogous proof for continuous time evolution is more subtle and is described in Balasubramanian:2022tpr.